Suppose you want to test the claim that μ ≠3.5. Given a sample size of n = 47 and a level of significance of α = 0.10, when should you reject H0 ?

A..Reject H0 if the standardized test statistic is greater than 1.679 or less than -1.679.

B.Reject H0 if the standardized test statistic is greater than 1.96 or less than -1.96

C.Reject H0 if the standardized test statistic is greater than 2.33 or less than -2.33

D.Reject H0 if the standardized test statistic is greater than 2.575 or less than -2.575.
Answer #1

Solution :

n = 47

df = 46

= 0.10

/ 2 = 0.10 / 2 = 0.05

t/2 = Z 0.05,46 = +/-1.679

A.Reject H0 if the standardized test statistic is greater than 1.679 or less than -1.679.